### v2

# Back from the dead: AT2019aalc as a candidate repeating TDE in an AGN

Abstract
Context. To date, three nuclear transients have been associated with high-energy neutrino events. These transients are generally thought to be powered by tidal disruptions of stars (TDEs) by massive black holes. However, AT2019aalc, hosted in a Seyfert-1 galaxy, was not yet classified due to a lack of multiwavelength observations. Interestingly, the source has re-brightened 4 years after its discovery.
Aims. We aim to classify the transient and explain the mechanism responsible for its second optical flare.
Methods. We conducted a multi-wavelength monitoring program (from radio to X-rays) of AT2019aalc during its re-brightening in 2023/2024.
Results. The observations revealed a uniquely bright UV counterpart and multiple X-ray flares during the second optical flaring episode of the transient. The second flare, similarly to the first one, is also accompanied by IR dust echo emission. A long-term radio flare is found with an inverted spectrum. Optical spectroscopic observations reveal the presence of Bowen Fluorescence lines and strong high-ionization coronal lines indicating an extreme level of ionization in the system.
Conclusions. The results suggest that the transient can be classified as a Bowen Fluorescence Flare (BFF), a relatively new sub-class of flaring active galactic nuclei (AGN). AT2019aalc can be also classified as an extreme coronal line emitter (ECLE). We found that, in addition to AT2019aalc, another BFF AT2021loi is spatially coincident with a high-energy neutrino event. The multi-wavelength properties of these transients suggest a possible connection between ECLEs, BFFs and TDEs in AGN.
Key Words.:
galaxies: active – galaxies: Seyfert – quasars: emission lines – individual: AT2019aalc – Neutrinos1 Introduction
Tidal disruption events (TDEs) are rare transient events, which occur when a star approaches a supermassive black hole (SMBH) and the tidal forces of the latter rip the star apart (Rees 1988). About half of the star’s material is accreted around the black hole, generating a luminous outburst across the electromagnetic spectrum. TDE optical light curves are characterized by quick and colorless evolving but also fast (on monthly timescales) decaying emission. Several TDEs were detected in X-rays with a very soft spectrum typically explained by late-time accretion disk formation (Saxton et al. 2020). In addition, a handful of TDEs produced detectable radio emission, explained by delayed111Note that the SMBH accretion rate does not peak on the same timescale as the mass fall-back rate. non-relativistic outflows (Horesh et al. 2021; Cendes et al. 2023), or only in instances by newly launched relativistic on-axis jets (Berger et al. 2012; Brown et al. 2015; Cenko et al. 2012; De Colle & Lu 2020; Yao et al. 2023a). Spectroscopically, TDEs are characterized by a strong blue continuum and broad emission lines ( km s-1) e.g., H and H. In a few cases, Bowen Fluorescence (BF: Bowen 1934, 1935) lines e.g., N iii at appear (e.g., Blagorodnova et al. 2019; Charalampopoulos et al. 2022a).
TDEs generally result in one optical flare, however, in a few cases a re-brightening episode was detected, even years after the initial flare. These might be explained in the following way. Under special circumstances, a star on a grazing orbit might be partially disrupted and only a fraction of the stellar mass will be deposited onto the SMBH. The secondary disruption of a surviving core after it reaches the pericenter again results in re-brightening. Only a few partial TDEs or candidates are known to date. These are AT2020vdq (Somalwar et al. 2023), ASASSN 14ko (Payne et al. 2021), eRASSt J045650.3-203750 (Liu et al. 2023), RX J133157.6-324319.7 (Hampel et al. 2022), AT2018fyk (Wevers et al. 2019, 2023), AT2019avd (Chen et al. 2022), AT2022dbl (Lin et al. 2024) and AT2021aeuk (Sun et al. 2025).
Based on the observed properties of TDE host galaxies, French et al. (2020) estimated a TDE rate of per year per galaxy which is well borne out by recent data (Yao et al. 2023b). Importantly, a similar fraction of TDEs should take place in AGN hosts (e.g., Chan et al. 2019; Ryu et al. 2024). These transients are challenging to distinguish from regular AGN variability. Only a few candidate TDEs taken place in AGN (in the following: TDE-AGN) were studied earlier (e.g., PS16dtm - a TDE in a narrow-line Seyfert-1 (NLSy1) galaxy (Blanchard et al. 2017) or the multiple soft X-ray flares in IC 3599 (Campana et al. 2015) and GSN 069 (Shu et al. 2018)). In contrast with TDEs seen in earlier quiescent galaxies, these transients are characterized by slowly decaying optical emission and bumps tend to appear after the peak.
The family tree of AGN-related transients became much broader in the past years (e.g., Frederick et al. 2021). In addition to TDEs, flaring NLSy1s, changing-look AGN (CLAGN) and different types of SNe were classified as typical nuclear transients. Recently, Trakhtenbrot et al. (2019) identified a new class of spectroscopically unique flares from accreting SMBHs. In , the transient events AT2017bgt, F01004-2237 and OGLE17aaj have been identified as the first members of the class based on their unique optical spectroscopic features. Later, the peculiar X-ray transient AT2019avd (Trakhtenbrot et al. 2020) was found with remarkable common properties, however, its classification as an AT2017bgt-like transient is unclear. A more promising case is AT2021loi (Makrygianni et al. 2023), a UV-bright transient event with very similar multi-wavelength properties to the instances studied by Trakhtenbrot et al. (2019). These events show unobscured AGN-like optical spectra with significant and persistent BF lines, extremely strong UV emission and slowly, on yearly scales, decaying optical emission. These AGN are characterized by significantly and longer-term intensified accretion. Hereafter, we will refer to this new class of SMBH flares as Bowen Fluorescence Flares (BFFs), following Makrygianni et al. (2023). Using the Zwicky Transient Facility (ZTF, Masci et al. 2019) Public Survey data, Dgany et al. (2023) found only one BFF candidate (AT2021seu) among spectroscopically classified transients, suggesting that these flares are rare. Two other candidates, namely AT2022fpx, (Koljonen et al. 2024) and AT2020afhd (Arcavi et al. 2024) were found later.
In recent years, two nuclear transients were found to be spatially coincident with neutrino events detected by the IceCube Neutrino Observatory based on ZTF follow-up campaigns of IceCube neutrino alerts (Stein et al. 2023). The TDE AT2019dsg is likely associated with the neutrino, IC191001A (Stein et al. 2021) and a further promising association, between the candidate TDE AT2019fdr and IC200530A (Reusch et al. 2022) points to TDEs as a potential new class of sources of the extragalactic neutrinos. Both sources were identified as having luminous mid-infrared emission, and a search for similar flares (van Velzen et al. 2024) led to the identification of the transient event AT2019aalc (a.k.a. ZTF19aaejtoy) with the neutrino alert IC191119A. In addition, two candidate obscured TDEs (Jiang et al. 2023) and the extremely energetic candidate TDE AT2021lwx (Yuan et al. 2024) were associated with high-energy neutrinos. A recently claimed association between a neutrino flare detected by the IceCube Observatory and an X-ray flare from the TDE ATLAS17jrp suggests that the X-ray emission of these sources play a crucial rule in producing neutrinos (Li et al. 2024).
In this paper, we focus on the nuclear transient event and neutrino source candidate AT2019aalc especially on its significant optical re-brightening that has started years after its initial flare (Veres et al. 2023).
The paper is organized as follows. We explain the discovery of the transient event and give an insight about its host galaxy in Sect. 2, summarize our observing campaign and the data reduction in Sect. 3, present the results of the observations in Sect. 4 and discuss them in Sect. 5 and finally give a summary and conclude in Sect. 6. Throughout the paper, we adopt a flat CDM cosmological model with parameters , , and . In this model (Wright 2006), mas angular size corresponds to pc projected linear size at the source redshift (Ahn et al. 2012). This redshift corresponds to a luminosity distance of Mpc. All magnitudes are given in AB system (Oke 1974).
2 The transient event: AT2019aalc
An analysis of the infrared properties of AT2019dsg and AT2019fdr revealed a strong reverberation signal i.e. the infrared dust echo, known as reprocessed emission of optical/UV photons in the infrared by hot dust around the SMBH, at a distance of pc (Jiang et al. 2016; Dou et al. 2016; Lu et al. 2016; van Velzen et al. 2021b; Jiang et al. 2021; van Velzen et al. 2024). In addition to the exceptionally high infrared luminosities of these transients, the neutrino detection times appear to be temporally consistent with the infrared luminosity peaks (Stein et al. 2021; Reusch et al. 2022). This discovery motivated a systematic search for neutrino emission using an extended sample of black hole flares (van Velzen et al. 2024). This archival search revealed the nuclear transient AT2019aalc with a powerful optical flare and dust echo emission in spatial coincidence with the high-energy neutrino event, IC191119A (IceCube Collaboration 2019) detected by the IceCube Observatory.
The host galaxy of AT2019aalc (2MASX J15241665+0451192) is a barred spiral (SBa) galaxy hosting an active nucleus (Oh et al. 2013). Based on its Sloan Digital Sky Survey (SDSS: Abazajian et al. 2009) spectrum it was further classified as a broad-line Seyfert-1 galaxy (Trump et al. 2013; Liu et al. 2019). The galaxy was detected in archival radio sky surveys e.g., the Very Large Array Faint Images of the Radio Sky at Twenty-cm (VLA FIRST: Condon et al. 1998). Fig. 1 shows the Pan-STARRS (Chambers et al. 2016) i-band optical image of the galaxy together with the contours of the FIRST radio data (at GHz) and the position of the X-ray counterpart detected by Swift/XRT in . The galaxy is not present in any X-ray catalogs prior to the optical discovery of the transient in .
In addition to the significant IR echo, AT2019aalc shares multi-wavelength characteristics with the two other neutrino candidate transients: a soft X-ray spectrum and transient radio emission (van Velzen et al. 2024). Moreover, similarly to the other two instances, the infrared peak of the transient event is temporarily coincident with the detection time of the likely associated neutrino event. Notably, Winter & Lunardini (2023) estimated the highest neutrino fluence for AT2019aalc among the studied candidates, due to its high estimated SMBH mass (, van Velzen et al. 2024) and low redshift. Since the proposed association with the high-energy neutrino event was announced roughly years after the discovery of the transient event, the source was not at the focus of attention. Thus, the lack of spectroscopic monitoring during the first optical flare hindered a clear classification of the event. Nevertheless, out of the more than AGN detected by the ZTF, less than show similarly rapid and large outbursts (Reusch et al. 2022; van Velzen et al. 2024) suggesting a very unusual transient. Interestingly, roughly years after the first flare, a significant optical re-brightening started in mid-May 2023 (Veres et al. 2023), illustrated by the long-term ZTF difference light curve shown in Fig. 2.
3 Observations and data reduction
3.1 Optical/UV
3.1.1 ZTF
ZTF forced photometry at the location of AT2019aalc was obtained following the recommendations outlined in Masci et al. (2023). We corrected for extinction by adopting a galactic extinction of (Schlafly & Finkbeiner 2011). As shown in Fig. 2, the forced photometry yields a negative flux in the period between the two flares (and also just before the first flare). This implies that during these epochs, the flux decreased below the mean flux of the reference images. For the -band these were obtained between March and May, 2018. While the baseline used to obtain the difference flux of the -band is based on images obtained between March and September, 2018.
3.1.2 UVOT
The Neil Gehrels Swift Observatory (Arnaud 1996a) performed Target of Opportunity (ToO) observations (project codes: 18972, 19188, 19993, 20997 and 21692, Proposers: Reusch, Veres & Sniegowska) on AT2019aalc from 2023-06-23 to 2025-01-02 with a total exposure time of ks. The Ultraviolet and Optical Telescope (UVOT, Roming et al. 2005) onboard the satellite observed the target field with the following four filters in each epoch: UW2 (central wavelength, 1928 ), UVM2 (2246 ), UW1 (2600 ) and U (3465 ). We analyzed the UVOT images using the Swift UVOT tools included in the heasoft 6.30 software package. We measured the flux using uvotsource, applying a circular aperature of arcseconds (which captures nearly all the flux of the host galaxy). The background was measured using four 6 arcsecond regions in four quadrants away from the target source. To estimate the UV flux before the flare, we use GALEX (Martin et al. 2005) observations obtained in 2007 and 2011. We use the same 15” circular aperture that was applied to the Swift/UVOT observations and use the gPhoton software (Million et al. 2016) to extract the flux. Over the 4 year baseline of the GALEX observations, the NUV magnitudes change by about 0.3 mag. Following van Velzen et al. (2021a), we combine the mean GALEX flux with the SDSS (York et al. 2000) magnitudes and apply Prospector (Johnson & Leja 2017) to find the best-fit stellar population synthesis model (Conroy et al. 2009) that describes these data. From this model we estimate the baseline flux in the UVOT filters and this baseline is subtracted to obtain the difference flux. The baseline flux in the UVM2 band is 17.2 mag, implying an increase of 1.8 mag relative to the peak of the UVM2 light curve. This large increase implies that, at the peak of the flare, we are not sensitive to the value of the baseline flux. This is useful because the baseline of the ZTF forced photometry was obtained around 2018 (at the start of the ZTF survey), while the baseline that is used to obtain the UVOT difference flux is mainly determined by the GALEX observations that were obtained a decade earlier.
3.2 Infrared
The Palomar 200-inch (P200) near-infrared telescope equipped with the Wide Field Infrared Camera (WIRC, Wilson et al. 2003) observed AT2019aalc on 2023-07-05. The imaging was obtained in the J-, H,- and Ks-bands. The data was processed with the Swarp software (Bertin et al. 2002). We performed background subtraction using the photutils.background class and measured the flux using a circular aperture with a radius of pixels. The corresponding zero-point magnitudes for this aperture size were calculated using the automatic pipeline for each filter. The error bars are the computed Poisson and background error of the photometry222https://www.stsci.edu/hst/instrumentation/wfc3/data-analysis/photometric-calibration/ir-photometric-calibration and we also took into account the uncertainty of the derived zero-point magnitudes.
We obtained mid-infrared photometry from archival observations of the Wide-Field Survey Explorer (WISE, Wright et al. 2010), which was continuously monitoring the sky at (W1) and, due to the NEOWISE project (Mainzer et al. 2011), at (W2). It visits any point in the sky roughly every six months for about ten times. Using timewise (Necker & Mechbal 2024; Necker et al. 2024), we downloaded all available multi-epoch photometry data from the AllWISE data release and single-epoch photometry from the NEOWISE-R data releases and stacked the data per visit to obtain robust measurements. All selected single-exposure photometry data points are within arcsecond of the nucleus of the host. The resulting light curve spans almost years, starting in February until July 2024.
3.3 X-ray
The X-ray Telescope (XRT, Burrows et al. 2005) observations were obtained simultaneously with the UVOT observations. The observations were performed in photon counting mode. We performed the data reduction using the Space Science Data Center (SSDC) interactive archive in the following way. First, cleaned event-files were produced using the xrtpipeline and xrtexpomap tasks. The ximage tool allowed us to determine the source detection level and its position. In case of non-detections, upper limits for the count rate were derived and converted to flux using WebPIMMS333https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3pimms/w3pimms.pl and assuming the mean power-law index of the observations that resulted in detections. When our target source was detected above the level, we ran swxrtdas in order to extract the source and the background spectra. Source counts were retrieved from a circle with a radius of pixel ( arcseconds), while background counts were extracted using an annular region with pixel inner/outer radius. Finally, the grppha tool of the heasoft 6.32 software package was used to perform background subtraction using the source and background spectra. We binned each spectrum to have a minimum of counts per spectral bin.
We fit the individual X-ray spectra using the XSPEC fitting environment (Arnaud 1996b). First, we have determined the Galactic neutral hydrogen column density, , based on the HI4PI survey (HI4PI Collaboration et al. 2016). The spectrum was then fitted using Cash-statistic (C-stat, Cash 1976). To find the best-fitting model, we first fitted the stacked spectrum with a absorbed power-law (Tbabs*powerlaw in XSPEC) model as the last XMM data of AT2019aalc before our monitoring program was best-fit with a power-law model (Pasham 2023). When fitting the model, the galactic column density was fixed whilst the power-law index and the normalization were free to vary. The resultant fit quality is C-Stat for degrees of freedom. We than added a blackbody model component, to take into account a possible soft excess. The blackbody temperature and the normalization of the new component were also free to vary and we fit again which resulted in a better fit with C-Stat for degrees of freedom. We implemented the likelihood ratio test (lrt in XSPEC) to estimate the statistical significance of the inclusion of the blackbody component. After running simulations, the test resulted only in smaller likelihood differences than the observed one clearly implying that the difference is unlikely to be due to random fluctuations. Therefore, accounting for the soft X-ray excess results in a significantly better fit.
3.4 Radio
3.4.1 VLBI observations
We observed AT2019aalc with the European VLBI Network (EVN) and the enhanced Multi-Element Remotely Linked Interferometer Network (e-MERLIN, Garrington et al. 2004). These observations were carried out at GHz (L-band) on 2023 June 09 (project code: EV027, PI: Veres). The observation time was hr. The experiment was performed in phase-referencing mode with -min duty cycles (with spending min on the target during each cycle) which resulted in an on-source time of h including fringe-finder scans and slewing. The data were recorded in left and right circular polarizations with a data rate of Gbit s-1. The correlation was done at the EVN Data Processor (Keimpema et al. 2015) at the Joint Institute for VLBI European Research Infrastructure Consortium (Dwingeloo, The Netherlands) with an averaging time of s. The Compact Symmetric Object (CSO) 15210430 (a.k.a. 4C 04.51) at an angular distance of was observed as phase-reference calibrator.
The data were calibrated with the U.S. National Radio Astronomy Observatory Astronomical Image Processing System (AIPS, Greisen 1990) software package following the standard procedures including time flagging, amplitude calibration, correction for the dispersive ionospheric delays and global fringe-fitting of the calibrator sources. We imaged the phase-calibrator with the Difmap software (Shepherd et al. 1994), using the hybrid mapping technique (Högbom 1974). Using the clean component (CC) model we built up in Difmap, we repeated the fringe-fitting in AIPS for the phase-reference calibrator source to correct its structure. Finally, we interpolated the solutions of the last fringe-fitting to the data of our target source. We imaged the target source using the hybrid mapping technique in Difmap. At the end of the procedure, several steps of amplitude and phase self-calibration were performed. A simple elliptical Gaussian modelfit component (Pearson 1995) was fitted directly to the visibility data. Finally, to decrease the noise level, we performed clean iterations in the residual image with a small loop gain of .
3.4.2 ATCA monitoring
The Australia Telescope Compact Array (ATCA)444https://csiropedia.csiro.au/australia-telescope-compact-array/, located at Narrabri, New South Wales, Australia, was used to monitor AT2019aalc. The observations were conducted between 2023-06-28 and 2024-05-04, utilizing the array in the 6km (6A and 6D) and (H168) configuration. The observations were conducted at frequencies of GHz, GHz, GHz, GHz and GHz. The flux calibrator was the standard ATCA primary calibrator, 1934–638, and 1548+056 is used as phase calibrator. The observations are summarized in Table LABEL:table:atca.
The data reduction procedures were based on the Multichannel Image Reconstruction, Imaging Analysis and Display (miriad: Sault et al. 1995). The data reduction procedure was conducted in accordance with standard protocol. In order to mitigate the impact of radio frequency interference (RFI) during flux and phase calibration, the interactive flagging tool bflag was utilised. Furthermore, automated flagging routines pgflag were applied to address interference for the source. Any corrupted data identified during calibration was manually flagged.
Due to the limited ()-coverage and the low elevation of the source resulting in a very elongated beamshape, we were not able to perform self-calibration and cleaning. Fortunately, AT2019aalc was always dominating the flux within the resolution elements of our short observations. Therefore, we imaged the region around our source for each observation using an oversampling of the synthesised beam by a factor of at least 10. Under the assumption that AT2019aalc is always a point-source, we can take the highest pixel value in the image as our measured flux. Due to the uncertainties this method introduced we used a plain % flux error.
Date | Exp. time / min | Configuration | Central Freq. / GHz | Bandwidth / GHz |
28.06.2023 | 330 / 180 | 6D | 2.1 / 7.0 | 1 / 4 |
30.06.2023 | 90 / 180 | 6D | 2.1 / 7.0 | 1 / 4 |
08.07.2023 | 120 / 120 | 6D | 2.1 / 7.0 | 1 / 4 |
19.07.2023 | 210 | 6D | 2.1 | 1 |
09.08.2023 | 120 / 150 | 6D | 2.1 / 7.0 | 1 / 4 |
19.10.2023 | 180 / 180 | H168 | 2.1 / 7.0 | 1 / 4 |
09.02.2024 | 60 / 60 / 60 | 6A | 2.1 / 7.0 | 1 / 4 / 4 |
29.03.2024 | 60 / 60 / 60 | 6A | 2.1 / 7.0 / 17.0 | 1 / 4 / 4 |
3.5 Spectroscopic observations
The Low Resolution Imaging Spectrograph (LRIS: Oke et al. 1995) mounted on the Keck-I telescope at Maunakea measured the spectrum of AT2019aalc, on 2021-07-06 (PI: Kulkarni). Later, during the second optical flare, we obtained more optical spectra of AT2019aalc using the following instruments: LRIS (PI: Kulkarni), DeVeny Spectrograph mounted on the Lowell Discovery Telescope (LDT) in Arizona (PI: Hammerstein) and the Double Beam Spectrograph (DBSP, Oke & Gunn 1982) mounted on the inch telescope at Palomar Observatory, California (PIs: Kulkarni and Kasliwal). The spectra were reduced following the standard procedures using the automated reduction pipeline of the LRIS (Perley 2019), the software package PypeIt (Prochaska et al. 2020) for DeVeny, and the software packages DBSP DRT (Mandigo-Stoba et al. 2022) and PypeIt (Prochaska et al. 2020) for the DBSP. Details of these observations of AT2019aalc are summarized in Table 2 while the spectra are shown in Fig. 17.
To determine the type of AGN prior the flares, we performed a stellar population synthesis fit of the pre-flare SDSS spectrum with (pPXF: Cappellari & Emsellem 2004; Cappellari 2017, 2023) and the newest GALAXEV (Bruzual & Charlot 2003) models (CB2019), assuming a Kroupa (2001) initial mass function (IMF) and an upper mass limit of , as templates (see Fig. 18). Beforehand, the spectrum was corrected for Milky Way extinction using the dust reddening maps by Schlafly & Finkbeiner (2011) and Cardelli et al. (1989) extinction curve. Furthermore, it was redshift corrected and logarithmically rebinned. We included several emission lines to the fit and assumed four kinematic components: One for all stellar templates, one for all forbidden lines, one for all allowed emission lines, and a last one for a second, broad component of all Balmer lines up to . Additionally, we included two separate dust components, each one with the extinction curve of Cardelli et al. (1989): a first one for stellar, and a second one for nebular emission. We therefore fixed the Balmer line ratio and limited doublets of the preinstalled lines to their theoretical ratio using the keywords tie_balmer and limit_doublets of pPXF. However, we only fixed the ratio for the strongest lines of the Balmer series up to to avoid possible contamination of by the [Ne iii] line. For the fit with pPXF we included additive and multiplicative polynomials up to tenth order (degree=10, mdegree=10) to model the AGN continuum. In order to correct for intrinsic dispersion caused by the spectrograph, we assume a resolving power of at and at .555https://www.sdss4.org/dr17/spectro/spectro_basics/ Linear interpolation is used to calculate values for wavelengths in between.
The newly observed spectra were fitted using the galaxy spectrum fitting tool PyQSOFit (Guo et al. 2018; Shen et al. 2019; Ren et al. 2024). The continuum of each spectrum was modeled with a combination of the following components: a power-law to represent the AGN continuum, polynomials to capture any residual smooth variations, and an Fe II template for the iron emission complex. Emission lines in the H and H complexes were modeled using a combination of broad and narrow Gaussian profiles to represent different kinematic components. The high-ionization coronal lines of interest were fitted with additional Gaussian profiles. The H complex shows a strong double-peaked feature around . This region is known as Bowen Fluorescence complex. It was fitted with a set of broad Gaussian components. We computed parameter uncertainties using MC resampling. Examples of the fitted spectra and the inferred emission line parameters are given in Fig. 19 and in Table 3, respectively.
4 Results
4.1 Optical/UV
The optical and UV light curves of AT2019aalc are shown in the bottom panel of Fig. 3. The transient was first detected by the ZTF in January with an -band magnitude of mag. AT2019aalc brightened to a peak magnitude of () reached on 2019-06-18 on a timescale of days. The second flare started mid-May , approximately years after the initial flare (Veres et al. 2023). The second flare evolved very similarly to the original one. From the beginning of the second flare, a mag rise was seen up to the peak of the flare on 2023-07-13 with mag (), on a timescale of days. The second flare peaked at roughly times the peak luminosity of the first flare. The peak of the second flare was followed by a short plateau phase that lasted for around days. Both flares declined very slowly, on timescales of years. The first flare decayed with a power-law index of while the second with in the -band. The decaying phases are not monotonic, as bumps tend to appear. During the second optical flare, these bumps appear four times: end of December , mid-March , end of May and mid-August (see Figure 3). Both optical flares brightened with constant color () and showed reddening when decaying.
Prior to the discovery of AT2019aalc, the host galaxy of the transient was monitored by several optical surveys. To study the pre-discovery optical variability of AT2019aalc, we performed aperture photometry of the All-Sky Automated Survey for Supernovae (ASAS-SN: Shappee et al. 2014; Kochanek et al. 2017) and the Catalina Real-Time Transient Survey (CRTS: Djorgovski et al. 2011). In addition, Palomar Transient Factory (PTF Law et al. 2009) magnitudes were extracted from the PTF Lightcurve Table666https://irsa.ipac.caltech.edu. No significant optical variability is visible between April and the discovery of the transient in January by ZTF (see Fig. 4).
The UV monitoring of the source started on 2023-06-23 around days before the optical peak of the second flare. The U-band peak is coincident with the optical peak while the shorter wavelength bands peaked weeks later. We note that right after the -band peak, a short plateau phase started in optical and UV. The transient is most luminous at the shortest wavelength ranges covered by UVOT. The UV emission peaks around erg/s in the filter UW2. After the peak, the UW2U color shows reddening. Towards the end of the UVOT monitoring, the measured UM2 magnitudes decreased back to the level of the baseline GALEX-NUV magnitudes measured between and .
Comparing the peak UVM2 magnitude of the UVOT monitoring to the GALEX-DR5 (Bianchi et al. 2011) NUV magnitude implies that the UV luminosity increased by a factor of with respect to the archival flux. The amplitude of this flare is higher than typically seen for AGN. Only out of the AGN presented in the GALEX Time Domain Survey (Gezari et al. 2013) have larger amplitude NUV variability.
4.2 Infrared
We detected AT2019aalc in all of the three bands observed with the WIRC. The observed magnitudes are plotted in the top panel of Fig. 3. We compare these to pre-flare magnitudes presented in the 2MASS All-Sky Catalog of Point Sources (Cutri et al. 2003). The observed magnitudes in April are higher with mag in J-, and Ks-bands and with mag in H-band indicating the clear brightening of the IR counterpart with respect to archival data.
The top panel of Fig. 3 shows the light curve starting just before the first optical flare. The distinct first dust echo flare that peaks around six months after the first optical peak. The rise of the dust echo of the second flare becomes also visible in the latest WISE data releases. The light curve varies by around mag in W1 and mag in W2 before the dust echo flare, which is consistent with statistical AGN variability (Berk et al. 2004). In contrast, the dust echo presents an increase of mag in W1 and mag in W2 after the first optical flare. Following the second optical flare, the IR light curve shows a brightening of around mag in both bands and is increasing when the optical flare is already decaying. Because the brightness is significantly above the limiting magnitude where Eddington bias becomes an issue for stacked single-exposure photometry (Necker et al. 2024), we can estimate the pre-flare mid-IR color using the median of the light curve prior to the dust echo flare with . Although there is clear AGN activity prior to the dust echo flare, this is still below the widely used AGN identification cut of (Stern et al. 2012). However, according to the reliability of WISE selection as a function of a simple W1W2 color selection (Stern et al. 2012), the color of AT2019aalc suggests a reliability of only %. With the onset of the dust echo, the IR emission becomes dominated by the heated dust. The change from bluer to redder colors comes from cooling of the dust. The re-brightening of the second dust echo is consequently accompanied by another color change towards the blue shown in Fig.5.
Notably, AT2019aalc is not part of the Flaires sample (Necker et al. 2024), a list of dust-echo-like infrared flares interpreted as extreme AGN accretion events or TDEs. This is because the broad application of the Flaires pipeline necessarily included a strict cut on extraneous variability that AT2019aalc did not pass. The authors did note however that a desirable relaxation of this criterion would include AT2019aalc. The source is, however, part of the dust echo sample of van Velzen et al. (2024) and, notably has the highest dust echo flux of all ZTF transients.
4.3 X-ray
The X-ray luminosity evolution is shown in the bottom panel of Fig. 3 while the count rate variability is plotted in the top panel of Fig. 6. The first flaring episode peaked at in the end of June (see Fig. 3.) This flare peaked weeks before the optical peak. The count rate dropped by a factor of within a timescale of days. Later, the X-ray emission shows limited variability for several months. A hint of increased emission was detected in mid-February which was also followed by a bump in the optical and UV light curves around weeks later. In the end of April , a rapid and extreme flaring episode started in X-rays that lasted less than a month but reached a luminosity of erg/s which is almost a magnitude larger compared to the peak of the first X-ray flaring episode. Interestingly, this X-ray flare was also accompanied by a bump in the optical weeks later. A similarly luminous and rapid flare was detected in August and once more, accompanied by an optical bump weeks later. Around days passed between the second and third and the third and fourth X-ray peaks.
In Fig. 7 we show the source spectrum derived from the observing campaign. The spectral index of the stacked spectrum fitted with a power-lawblackbody model is while the blackbody temperature is eV. These imply an overall unabsorbed luminosity value of L erg/s. The mean count rate of the observations is cts s-1. The photon index variability of AT2019aalc is presented in the bottom panel of Fig. 6.
AT2019aalc (= SRGe J152416.7045118) was observed by the eROSITA (Sunyaev et al. 2021) soft X-ray telescope four times starting from 2020-02-02 (roughly half a year after the peak of the first optical flare) with a months cadence. The X-ray light curve reached a plateau between August and January with a flux of F erg cm-2s-1 in the energy range of keV. The source had a soft thermal spectrum described with a blackbody temperature of eV (van Velzen et al. 2024). One observation with XMM-Newton Observatory (Jansen et al. 2001) was performed on 2021-02-21 during a period of optical quiescence. The best-fit power-law index and the observed keV flux of this observation were and F erg cm-2s-1 (68 confidence), respectively (Pasham 2023).
4.4 Radio
4.4.1 Temporal evolution
The pre-flare system was detected in the NRAO VLA Sky Survey (NVSS, White et al. 1997) in February 1995 and in the FIRST survey in February 2000 (both at GHz) with peak flux densities of mJy and mJy respectively, suggesting no substantial variability. These flux densities correspond to a monochromatic radio luminosity of W Hz-1 measured at GHz which is typical for Seyfert galaxies (Meurs & Wilson 1984).
Later, the transient’s sky location was observed in the Karl G. Jansky Very Large Array Sky Survey (VLASS, Lacy et al. 2020). The VLASS observes the entire sky visible to the VLA at GHz wide band three times. Our source was detected with peak flux density values of mJy/beam (on 2019-03-14), mJy/beam (on 2021-11-06) and mJy/beam (on 2024-07-20), corresponding to 3 GHz radio luminosities () of , and , respectively. This factor of increase in flux from the first VLASS epoch (three months before the first optical peak) to the second VLASS epoch (2 years post-peak) is statistically significant (at the 8-level, as estimated from the rms in the VLASS “Quick Look” images, van Velzen et al. 2024).
In addition, we found a radio source at the position of AT2019aalc in the first data release of the Rapid ASKAP Continuum Survey (RACS: McConnell et al. 2020). The source was observed in two different fields and, consequently, at two different epochs. The observed peak flux densities of mJy/beam (on 2019-04-24) and mJy/beam (on 2020-04-30) correspond to a MHz radio luminosity of and indicate that the radio flare seen in the VLASS data started at least days after the peak of the first optical flare, sometime between April 2020 and November 2021.
The radio luminosity in the beginning of our ATCA monitoring was at the same level ( ) when the source was observed for the second time in the VLASS in April 2021. Since the radio brightness does not show rapid variability during the months of our ATCA monitoring, we can reasonably assume that the sharp increasing resulted in a long-lasting plateau. We present the ATCA radio multi-frequency light curves of AT2019aalc together with the archival observations in the middle panel of Fig. 3.
4.4.2 The radio spectrum
To further characterize the radio emission of AT2019aalc, we derived the radio spectral index of the radio source in each epoch of our ATCA monitoring. The spectral index between GHz and GHz (, defined as ) does not show any obvious signs of variability on a timescale of months. We give an average spectral index of . We calculated a pre-flare spectral index using the VLASS epoch and RACS epoch flux densities . The spectral index we derived between GHz and GHz this way is .
Our high-frequency K-band observations centered at GHz, however, do not follow the steep power-law and suggest a turnover between GHz and GHz.
4.4.3 VLBI detection
The naturally-weighted VLBI map of AT2019aalc is presented in Fig. 9. We detected a single radio-emitting feature with a signal-to-noise ratio (SNR) of . We obtained the following parameters of the fitted component: a total flux density of mJy and a size of mas at a position angle of . We calculated the uncertainties of the flux density and size parameters following the formulas of Fomalont (1999). The minimum resolvable size () of a Gaussian component fitted to naturally weighted VLBI data was calculated following Kovalev et al. (2005). This yields mas implying that the fitted Gaussian component is resolved.
We calculated the brightness temperature of the radio-emitting feature found at the position of AT2019aalc using the following equation (Condon et al. 1982; Ulvestad et al. 2005):
where is the flux density of the fitted Gaussian component measured in Jy, and are the major and minor axes full width at half maximum (FWHM) of the fitted component in mas, and is the observing frequency in GHz. The obtained brightness temperature is K at GHz. The brightness temperature is an effective parameter that is commonly used in radio astronomy to describe the physical properties of emitting material in astrophysical objects (e.g., Lobanov 2015).
4.5 Spectroscopic results
We present the archival and the newly obtained optical spectra of AT2019aalc in Fig. 17 with marking the notable emission lines. In the following, we highlight the remarkable line detection of the pre-flare spectrum observed before the first optical flare, the one observed between the two flares (post-flare-1 spectrum) and five spectra that were taken during the second flare (over-flare-2 spectra).
4.5.1 Pre-flare spectrum
A spectrum of the pre-flare system was taken in April by the SDSS spectrograph. We present this spectrum with the best fits for the stellar and gas components in Fig. 18.
We identify strong Balmer emission lines. We calculated FWHM(Balmer) kms-1 for the broad component of the Balmer lines in the pre-flare spectrum, which is typical for an unobscured, broad-line AGN. The line widths (FWHM(Balmer) km s-1) presented in the SDSS DR7 broad-line AGN catalog (Liu et al. 2019) for our source are consistent with this classification.
Several forbidden lines such as [O iii] and [O iii] , [S ii] , [S ii] and [N ii] were detected. These lines are produced in the narrow-line region (NLR) in AGN. We detected the [O ii] [O ii] line doublet, which is a potential indicator of AGN-driven outflows (e.g., Santoro et al. 2020).
The pre-flare spectrum also shows the detections of two high-ionization coronal lines ([Fe vii] and [Fe xiv]) which are further discussed below.
4.5.2 Post-flare-1 spectrum
The first LRIS spectrum was taken almost years after the first optical flare and years before the second one.
In this spectrum, prominent Balmer lines were detected, similarly to the SDSS spectrum. The spectrum exhibits the aforementioned forbidden lines detected in the archival spectrum.
We identify the Bowen Fluorescence line O iii and He ii .
The high ionization coronal lines [Fe vii], [Fe x], [Fe xiv] and [Fe vii] were detected. We identify the [Ne V] emission lines at and in the spectrum. Note that the pre-flare SDSS spectrum does not cover these wavelengths, therefore these lines are not necessarily related to the optical flares discussed here. These lines are prominent emission lines originating from the inner NLR excited by the AGN via photoionization or shocks and are sometimes considered as coronal lines (e.g., Cann et al. 2020).
4.5.3 Over-flare-2 spectra
We performed optical spectroscopic observations of AT2019aalc during the second flare. These are characterized by a blue continuum with several emission lines appearing in comparison to the spectra discussed above. We compare the second LRIS spectrum of AT2019aalc (covering the broadest wavelength range) with a composite SDSS spectrum based on Seyfert 1 galaxies published in Pol & Wadadekar (2017) ,in Fig. 10. This comparison shows similarities but also significant differences discussed below.
Over the second optical flare, strong Balmer lines were detected. Two lines of the series (H and H) were detected only in the over-flare-2 spectra.
Over the second optical flare of AT2019aalc, we detected the Bowen Fluorescence line complex N iii +He ii . This complex was not detected in the earlier spectra. The relative strength of the BF complex to H reached a maximum on 2023-07-11, very close to the optical continuum peak (2023-07-13). As we detected the Bowen line N iii even days after the optical peak of the second flare, we conclude that it persists over long timescales. We also find hints of O iii which is also known as a Bowen line (Selvelli et al. 2007), however, due to a possible blending from Fe vii, its detection is questionable. Nevertheless, this Bowen line is created through the same channel (O1) as the above discussed O iii which supports the hint of detection. O iii was detected in some metal-rich TDEs before (Leloudas et al. 2019) such as AT2019dsg (Cannizzaro et al. 2021).
We detect the high-ionization coronal lines seen in the post-flare-1 LRIS spectrum over the second flare in each spectrum. Furthermore, a new coronal line appears: (Fe xi (ionization potential: eV). We also detected the [Ne V] emission lines at and in the second LRIS spectrum, as well as in the post-flare-1 spectrum.
5 Discussion
5.1 Spectroscopic properties
5.1.1 Balmer lines
AT2019aalc exhibits prominent Balmer lines in each spectrum. In addition, a more complex structure of these lines appears in the late-time over-flare-2 spectrum. The Balmer series lines here exhibit a redward wing which, however, in some cases, might result from blending with other lines, such as H and the BF line N,iii . Interestingly, the post-flare-I spectrum shows similar structure of the Balmer line series.
5.1.2 Bowen fluorescence lines
Wavelength coincidences between emission lines (”line fluorescence”) can be an important source of radiative excitation. The Bowen Fluorescence mechanism in astrophysical environments is a good indicator of absorbed EUVsoft X-ray flux (at wavelengths shorter than the He ii Lyman limit of eV) as these high-energy photons are converted into the He ii Lyman emission required for the excitation of the O iii and N iii lines in the optical and NUV regimes (Netzer et al. 1985; Netzer 1990). This is consistent with the multiple soft X-ray flares and the extreme UV luminosities of AT2019aalc. The BF line O iii has been observed in the Bowen Fluorescence Flares AT2017bgt (Trakhtenbrot et al. 2019) and AT2021loi and only in a few Seyfert galaxies before (e.g., Malkan 1986; Schachter et al. 1990). However, we note that due to the atmospheric cutoff at these lines are difficult to detect. This line is typically weaker in regular Seyfert galaxies than in the case of AT2019aalc (see Fig. 10). The BF line N iii is not generally seen in AGN (e.g., Vanden Berk et al. 2001), however, it has been observed in several optical TDEs (Charalampopoulos et al. 2022b) and other SMBH-related transients: BFFs (Trakhtenbrot et al. 2019; Makrygianni et al. 2023), the peculiar transient event AT2019avd (Malyali et al. 2021) and a few flaring NLSy1s (Frederick et al. 2021).
The detection of a BF line almost years after the first optical flare of AT2019aalc suggests its long persistence, similar to the BFFs studied by Trakhtenbrot et al. (2019). In the case of AT2017bgt, this line was still detected days after the discovery of the transient. The Bowen features in F01004-2237 were detected even years after the initial flare (Tadhunter et al. 2021). We compare the over-flare-2 high-resolution LRIS spectrum of AT2019aalc to the spectra of BFFs, a TDE-Bowen, and a repeating partial TDE in Fig. 11 around the N iii He ii complex. We further compare the second LRIS spectrum of AT2019aalc with the LRIS spectrum of AT2021loi in Fig. 12 zooming into the O iii He ii region.
The He ii ionized helium line is presented in the spectrum taken between the optical flares.. This emission line is commonly present in AGN and known as a recombination line, tracing ionized gas in the vicinity of the central SMBH responsible for the photoionization (e.g. Wang & Kron 2020). This line is known to be created by photoionization due to (soft) X-ray photons with a rough correspondence of He ii photon emitted for each keV X-ray photon (Pakull & Angebault 1986; Schaerer et al. 2019; Cannizzaro et al. 2021). In the spectra of AT2019aalc, He ii is detected together with strong N iii and O iii BF lines and He ii , indicating its relation to the Bowen Fluorescence mechanism. Furthermore, we calculated a maximum line intensity ratio of which significantly exceeds the typical ratios estimated for AGN (earlier Vanden Berk et al. (2001) estimated ratios of while Shirazi & Brinchmann (2012) gives ). Our value is, however, more comparable with the line ratio estimated for the BFF AT2017bgt (, Trakhtenbrot et al. 2019).
5.1.3 Coronal lines
Based on a study of SDSS spectra, Wang et al. (2012) reported the discovery of Extreme Coronal Line Emitter (ECLE) galaxies which show extremely strong coronal lines. Three TDEs (AT2017gge, AT2019qiz and AT2022upj) were also found to exhibit coronal lines (Onori et al. 2022; Newsome et al. 2022; Short et al. 2023). The BFF AT2021loi (Makrygianni et al. 2023) and the spectacular transient event AT2019avd (Malyali et al. 2021) also exhibit strong coronal lines. The BFF F01004-2237 showed variable coronal lines on a timescale of years. Based on the publicly available spectra of AT2017bgt on the Transient Name Server777https://www.wis-tns.org/ (TNS) we see enhanced coronal line emission in this BFF as well.
Wang et al. (2012) state that the most possible origin of the ECLEs are remnants of an earlier TDE outburst which was later confirmed by Short et al. (2023) who revealed that the TDE AT2019qiz more likely follows the line ratio correlations of ECLEs instead of regular AGN. The coronal lines in AT2019qiz appeared around days after the optical peak, further supporting this scenario. Moreover, the estimated rate of coronal line emitters is consistent with the estimated rate of TDEs (Wang et al. 2012). Later, Callow et al. (2024) discovered 9 new ECLEs using SDSS spectra and refined the previous rate calculations. They found that the ECLE rates are below the TDE rates, however, can still be consistent with assuming that only to per cent of all TDEs produce variable coronal lines. According to Wang et al. (2012), the origin of the coronal lines in ECLEs can be explained with the liberation of Fe from dust grains that were destroyed by the TDE flare. A follow-up spectroscopic analysis of the original ECLE sample of Wang et al. (2012) revealed that the coronal line signatures did not recur in five out of the seven objects, supporting their transient light echo origin (Clark et al. 2024). The large dust echo of AT2019aalc is consistent with this picture. Furthermore, the TDEs with coronal line detection have more luminous IR dust echo compared to other optical TDEs. AT2019qiz is one of the few TDEs included in the accretion flare sample of van Velzen et al. (2024) due to its strong dust echo. Another coronal line-detected TDE AT2017gge is part of the mid-infrared (MIR) dust echo flare sample of Hinkle (2024) which contains ambiguous nuclear transients with high dust covering factors. The extreme dust echo seems to characterize the BFFs as well. The BFFs AT2021loi, AT2017bgt, OGLE17aaj and the candidate BFF AT2019avd are present in the above-mentioned MIR dust echo sample (Hinkle 2024) and of them are also part of the Flaires sample of dust-echo-like IR flares (Necker et al. 2024). The Bowen line transient AT2019pev shows strong high-ionization coronal lines (Frederick et al. 2021; Yu et al. 2022) and is also part of the MIR flare sample of Hinkle (2024). These findings suggest a connection not only between the ECLEs and TDEs, but also the BFFs. We performed modeling of the publicly available LRIS spectrum of AT2017bgt taken on 2017-02-23 using PyQSOFit, following the procedure described in Subsect. 3.5. Fig. 13 shows the [Fe x] versus the [O iii luminosities of the ECLEs studied by Wang et al. (2012), AT2019qiz (Short et al. 2023), two BFFs and AT2019aalc together with the SDSS Seyfert sample of Gelbord et al. (2009). This clearly implies an offset of the ECLEs, BFFs and AT2019aalc from the regular AGN sample.
5.2 Optical re-brightening
The long-term ZTF light curve of AT2019aalc shows two well-separated flares. Both flares evolved quickly and reached their peaks within months, followed by a slow decay on yearly timescales. Generally, TDEs are characterized by smoothly and rapidly declining light curves () as expected from mass fallback considerations (Rees 1988) without any prominent bumps in the case of full disruption. AT2019aalc in turn has a slowly decaying first flare with a power-law fit returning a power-law index of and a second flare that began years after the initial flare. The first systematically identified repeating partial TDE (AT2020vdq), however, shows not just a double-peaked long-term optical light curve, but its second flare is times more luminous compared to the first one (Somalwar et al. 2023). This implies an even more significant optical re-brightening than the one detected in AT2019aalc. The second flare of partial TDEs is expected to follow a steep declining with a power-law of . In the case of AT2020vdq, the second flare declined much slower (Somalwar et al. 2023) than this but still clearly faster than the second flare of AT2019aalc which declined with a power-law index of .
Interestingly, most of the sources of the small sample of classified BFFs show optical re-brightening or bumps. Fig. 14 shows the optical evolution of the three BFFs with detected re-brightening episodes together with AT2019aalc. The optical light curve of AT2017bgt shows a bump roughly days after its initial flare which is also clearly visible in the binned light curve published in Trakhtenbrot et al. (2019). Originally, F01004-2237 had undergone a luminous flare (Tadhunter et al. 2021) in 2010 and later Trakhtenbrot et al. (2019) classified it as a BFF. The binned light curve of F01004-2237 published in Makrygianni et al. (2023) shows a bump around a year after its peak. We investigated the recent activity of this source and found that it significantly re-brightened in based on its ATLAS monitoring. The recent flare was more luminous than the first one and declined faster. In addition to these two BFFs, AT2021loi shows a bump when declining, roughly a year after the first flare reached its peak. Makrygianni et al. (2023) suggests that optical re-brightening or bumps seen when decaying might be common features of BFFs. Only one transient classified as a BFF, OGLE17aaj, has not shown signs of optical re-brightening or bumps in its light curve to date. The first flare of the studied BFFs decay on yearly timescales, much slower than regular TDE flares but more similar to AT2019aalc. However, the BFFs with re-brightening episodes tend to have a faster decaying second flare/bump in comparison with their initial flare, while in the case of AT2019aalc the second flare decays slower. The BFFs classified so far tend to have remarkable variances in their optical light curve evolution despite the similarities we see. These might be explained by different subclasses of BFFs, similarly to regular TDEs, where different subclasses were defined based on their light curve evolution (e.g., Charalampopoulos et al. 2023).
5.3 Radio characteristics
5.3.1 Possible explanations of the long-term radio flare
In this subsection we discuss the long-term radio variability of AT2019aalc. The source experienced a brightening of its flux density by a factor of on a timescale of less than years followed by a long-term plateau ( years) indicated by the archival radio survey data and our ATCA monitoring. Although radio variability on yearly timescales is a common phenomenon of Seyfert galaxies, the sharp, high-amplitude radio flare and especially the long-term plateau are generally not seen in intrinsically variable Seyfert galaxies (e.g., Mundell et al. 2009; Koay et al. 2016) and therefore require further investigation. The unusual radio spectrum with a high frequency turnover is also discussed below.
The brightness temperature of K at GHz significantly exceeds K, known as an upper limit for supernova remnants and He ii regions in star-forming galaxies (Condon 1992). Consequently, the radio emission we detected must be non-thermal and originate from physical processes associated to AGN-like (including TDEs) activity (see Bontempi et al. 2012 as an example how the estimated brightness temperatures help to characterize the radio emission of Seyfert galaxies). This way, we can also safely rule out thermal free-free emission from the accretion disk or the dusty torus (e.g., the case of NGC 1068, Gallimore et al. 1997) as the origin of the detected radio emission.
Advection-dominated accretion flows have also been proposed to explain the origin of the radio emission of low-luminosity AGN (e.g., Seyfert galaxies, Mahadevan 1997). Assuming that accretion flows dominate the radio emission of AT2019aalc, following Yi & Boughn (1998) we estimate an expected GHz radio luminosity of erg s-1, a magnitude of lower than the observed one, clearly indicating that these cannot explain the observed radio emission.
Magnetized coronal winds originating from the AGN accretion disk could produce the non-thermal radio emission of AT2019aalc. Laor & Behar (2008) found a correlation of studying the Palomar–Green (PG: Green et al. 1986) radio-quiet quasar sample (in good agreement with the correlation found for coronally active stars by Guedel & Benz (1993). Later, Behar et al. (2015) came to the same conclusion using a sample of seven radio-quiet Seyfert galaxies. For our source, based on the mean ATCA C-band and the extrapolated Swift/XRT keV luminosities, we estimate a ratio of suggesting that magnetized coronal winds alone cannot fully explain the observed radio emission of AT2019aalc. Moreover, coronal emission should be highly variable at radio wavelengths.
Out-flowing synchrotron emitting material induced by enhanced accretion are commonly detected in AGN and TDEs in the forms of non-relativistic outflows or relativistic jets (e.g., Paragi et al. 2017; Mohan et al. 2022; Kozák et al. 2024). The high brightness temperature of the EVN-detected radio feature indicates its radio core identification. The steep () radio spectrum indicate the presence of another, steep spectrum synchrotron-emitting component(s) in the system as radio cores are characterized by a flat radio spectrum. The spectrum may be the composition of the flat spectrum radio core and the additional, steep spectrum component(s). The high-frequency turnover around GHz supports the idea of a newly ejected radio-emitting component.
5.3.2 The radio spectrum
The turnover suggests the multiple component nature of the spectrum. Radio spectra with high-frequency excess have been observed in a few radio-quiet narrow-line Seyfert 1 galaxies (Berton et al. 2020). These sources are thought to be kinematically young AGN with small linear sizes or heavily absorbed AGN. The high-frequency peak might be explained with the kinematic age of a newly ejected component. These extremely young radio sources therefore peak at very high frequencies (e.g., Berton et al. 2020). The low-frequency peak is, however, more likely to be explained with star formation activity instead of AGN emission since the latter is absorbed due to synchrotron self-absorption (SSA) or free-free absorption (FFA).
However, in the case of AT2019aalc, the large-amplitude variability concluded from the VLASS observations and the EVN detection of a high brightness temperature ( K) feature at GHz clearly imply AGN-like activity over a star formation scenario at these frequencies and is incompatible with significant absorption caused by SSA or FFA. Therefore, our results indicate that the radio spectrum of AT2019aalc is strongly dominated by AGN activity below GHz as well, even if a contribution from star formation, especially at GHz, is probable.
A newly ejected component can explain the high-frequency excess whilst the low-frequency part is originating from optically thin synchrotron emission of the AGN. Thus, the overall spectrum represents a superposition of past and recent AGN activity, indicative of its recently enhanced, intermittent behavior. Järvelä et al. (2021) states that restarted AGN activity can explain at least some of the radio spectra studied by Berton et al. (2020). Intermittent activity is further supported by theoretical models highlighting that sources with high accretion rates are more prone to this kind of behavior (Czerny et al. 2009).
5.4 X-ray properties
The X-ray spectra of AGN at energies above keV have a power-law-like shape. These X-ray photons are believed to be triggered by Compton up-scattering of the accretion disk photons off hot electrons surrounding the disk, in a hot ( K) optically thin corona above the disk (e.g., Kammoun et al. 2015). Below keV energies many AGNs show an excess in their spectrum referred to as the soft X-ray excess. It can be well modeled by a blackbody model with a best-fit temperature in the range keV. The origin of the excess might be explained via Comptonized disk emission and the reflection of hard X-ray photons from the surface of the disk can account also for the excess (Done et al. 2012).
The X-ray spectrum of AT2019aalc exhibits a non-thermal component with a power-law index of and an unusually soft thermal component with a blackbody temperature of eV. Most of the BLSy1s typically have a harder X-ray spectrum with a power-law index of (e.g., Pons & Watson 2014). Our value is more commonly seen in sources characterized by enhanced accretion, e.g., flaring NLSy1s or TDEs which have typically soft X-ray spectra with (Boller et al. 1996; Saxton et al. 2020).
The soft X-ray flux of Seyfert galaxies was found to be well correlated with the observed fluxes in the coronal line [Fe x]. The coronal lines to X-ray flux ratio is log for both broad and narrow-line Seyfert 1 galaxies (Gelbord et al. 2009). For AT2019aalc we estimate values of log when optical spectroscopic and X-ray observations were performed nearly simultaneously (due to the rapid X-ray variability, we consider only observations taken on the same day). Thus, the soft X-ray photons excite the coronal lines continuously and not years later originating from a past TDE as suggested for ECLEs by Wang et al. (2012).
The reoccurring optical/UV bumps seen during the decaying of the main optical flares are likely driven by the soft X-ray flares which lead these outbursts only with a couple of weeks. The soft X-ray excess appears to be more dominant during these flares. Interestingly, the soft X-ray flares show hints of quasi periodicity of days. A few Repeating Nuclear Transients (RNTs) were discovered in the past few years: Quasi-Periodic Eruptions (QPEs) exhibit with soft X-ray flares on time-scales of hours to days while partial TDEs on timescales of months to years (e.g., Chakraborty et al. 2021; Quintin et al. 2023; Miniutti et al. 2023; Nicholl et al. 2024). We note that another BFF, AT2020afhd, shows hints of quasi-periodic soft X-ray flares according to its Swift/XRT monitoring. This source is included in the The Living Swift XRT Point Source Catalogue (LSXPS: Evans et al. 2023) as LSXPS J031335.6020905. The potential connection between BFFs and QPEs should be explored in future studies.
5.5 SED fitting
Following Reusch et al. (2022), we fitted two blackbody models in order to describe the SED of the transient using lmfit. These fitted curves to the optical/UV (’blue’ blackbody) and IR datapoints (’red’ blackbody) are shown in Fig. 15, together with a combined double blackbody fit. The red blackbody fit was introduced to account for the IR dust echo. These curves fit the SED around the optical peak of the investigated flare.
The continuum emission of TDEs is well described by a thermal blackbody model (Gezari 2021). The case of AT2019aalc appears to be more complex. The UV emission has a contribution from the Bowen Fluorescence mechanism and possibly also from Comptonization explaining the poor fit towards the blue. The excess seen in the WISE bands is possibly due to the non-simultaneous observation with respect to the optical/UV and WIRC observations as the IR light curve is peaked well after the optical one. The excess in the -band possibly has a contribution from the early appearance of some Fe lines in this range but such a significant excess cannot be fully explained with the coronal lines only.
The inferred blackbody temperature and radius values estimated for the ’blue’ blackbody fit are K and cm, respectively. The blackbody temperature is in the lower range of that is found for optical selected TDEs, however, an order of magnitude lower than the typical values of the X-ray detected TDEs studied by van Velzen et al. (2020) and Thomsen et al. (2022). On the other hand, the radius is smaller than calculated for optical selected TDEs but is more consistent with the X-ray detected ones. The ’red’ blackbody fit resulted in K and cm.
5.6 What type of transient is AT2019aalc?
AT2019aalc shares similarities with TDEs such as the soft X-ray spectrum, the dust echo and the optical flare(s) rising with constant color. The UV bright nature of the transient with a blue optical/UV peak color of NUV is another TDE-like feature and not typical for AGN and SNe (Gezari 2021). Bowen lines appear in the spectra of severeal TDEs even if these are typically weaker than in our case. The most remarkable difference is, however, the fading of the optical emission, which is slower than for any TDE identified so far. The reddening in color when the main optical flares are declining is also atypical for TDEs.
The unobscured AGN-like optical spectra, the presence of persistent and strong Bowen lines (O iii and N iii ), the extremely luminous UV flare, on yearly timescales decaying optical emission and the optical re-brightening altogether suggest a Bowen Fluorescence Flare classification of AT2019aalc. Furthermore, we classify AT2019aalc as an Extreme Coronal Line Emitter because of the detection of several strong high-ionization coronal lines.
The mechanism that powers the BFFs is still unclear. TDEs in AGN might play a role. In a single star case, the slowly decaying optical/UV emission might be explained by the dissipation of energy due to the continuous interaction between the stellar debris stream and the pre-existing accretion disk (similarly to the case of the TDE-AGN PS16dtm, Blanchard et al. 2017, and ASASSN-14ko, Huang et al. 2023) or the AGN’s radiation field. According to TDE-AGN simulations, stellar passages can induce substantial perturbations in the inner disk and leave it in a very perturbed state even for years (Chan et al. 2019). The debris-disk interaction results in enhanced accretion and multiple shocks (Chan et al. 2019), notably leading to X-ray emission (Ryu et al. 2024). Consistently within this picture, the soft X-ray flares of AT2019aalc suggest that the emission is linked to the accretion disk. These X-ray flares lead the optical/UV bumps by weeks, which are detected during the decay of the main flare. The soft X-ray emission is likely reprocessed in optical/UV by the dense environment explaining the optical/UV flares and the emergence of the Bowen fluorescence and coronal lines. The TDE-AGN scenario provides an explanation for one slowly decaying flare with recurring outbursts, however, AT2019aalc exhibits two distinct optical flares. TDE-AGN simulations also revealed that in a case of an initially partial disruption, a surviving core would generate similar heating of the inner disk to the first pericenter passage (Chan et al. 2019; Ryu et al. 2024). Notably, the second flare of AT2019aalc is more luminous than the first one, similarly to the cases of the partial TDE AT2020vdq (Somalwar et al. 2023) and the BFF F01004-2237. Numerical simulations demonstrated that the even more luminous second flare of a partial TDE might be explained in the following way: tidal energy injection leads to oscillations in the remnant star, and if the thermal energy cannot dissipate within one orbital period, the star expands, making it more susceptible to further tidal disruption and potentially resulting in brighter flares (Zhong 2024; Liu et al. 2024). Liu et al. (2024) claimed that these considerations provide a plausible explanation for the brighter second flare of the partial TDE AT2020vdq. A partial TDE scenario might explain the second optical flare of the other two BFFs (AT2017bgt and AT2021loi) with double-peaked optical flares as well.
Although the slowly decaying optical emission and the re-brightening seem to be common features of BFFs, the long-term light curves still show differences as well. The re-brightening episodes started at different timescales for the individual sources, and more BFFs need to be classified to better understand their unique optical light curves. Nevertheless, (partial) TDEs occurring in AGN can be a possible origin of BFFs. This would explain the TDE-like properties of these sources. The most remarkable differences, such as the slowly decaying optical emission and the continuum emission not being well described by blackbody models, may arise due to the more complex environment in which the TDE occurs, compared to the simpler conditions in quiescent galaxies (”TDEs in vacuum”). TDE-AGN simulations showed that the emergent spectrum of such a system may not be purely thermal as in regular TDEs (Chan et al. 2019). Furthermore, a possible connection between the high-ionization coronal lines (which are most probably related to TDE activity, Wang et al. 2012) and the BFFs (see Fig.13) is consistent with this picture. A known candidate partial TDE in an AGN is the case of the active galaxy IC 3599. This low-luminosity AGN exhibits periodic soft X-ray flares interpreted with a partially disrupted star, which results in reoccurring outbursts after each passage of the surviving core around the central black hole (Campana et al. 2015). Notably, IC 3599 has more prominent high-ionization coronal lines than most AGN (similarly to the coronal-line detected transients, see Fig. 13). In addition, IC 3599 is the only known AGN with fading coronal lines (Frederick et al. 2019).
5.7 Radio properties of BFFs
As we studied AT2019aalc in radio in detail, we compare the radio properties with that of the BFFs. Based on the radio observations explained in Makrygianni et al. (2023), AT2021loi has an inverted radio spectrum between and GHz while the source was not detected at GHz. Although no radio spectrum is available for AT2017bgt, we investigated the VLASS QL Images of this BFF to study its radio flux variability. We found a radio source at the optical position of AT2017bgt with flux densities of mJy (on 2019-05-21), mJy (on 2021-10-19) and mJy (on 2024-06-11) at a central frequency of GHz. The pre-flare system was detected in the FIRST (at GHz) in July 1998 with a flux density of mJy. The BFF AT2020afhd was detected with the VLA at GHz with mJy (Christy et al. 2024). If BFFs are explainable with TDE-AGN, the synchrotron emission is likely originating from outflow activity due to the enhanced accretion or, alternatively, from interaction between the unbound debris and the surrounding gas predicted by TDE-AGN simulations published in Chan et al. (2019). Radio observations of any newly discovered BFF will be essential to better understand these transients.
5.8 BFFs and high-energy neutrinos
AT2019aalc was associated with the high-energy neutrino event IC191119A (most probable neutrino energy: TeV) detected by the IceCube Observatory (IceCube Collaboration 2019; van Velzen et al. 2024). Motivated by the classification of AT2019aalc suggested by its multi-wavelength properties (Subsect. 5.6), we searched for spatial coincidences between the earlier classified BFFs and the neutrino alerts from the IceCube Event Catalog of Alert Tracks (ICECAT-1: Abbasi et al. 2023). We found that the BFF AT2021loi is located within the 90% rectangular uncertainty contour of the high-energy neutrino event IC230511 (most probable neutrino energy: TeV) detected in May (IceCube Collaboration 2023). In the full likelihood contours published in ICECAT-1, the source is just outside the 90% contour. We present the source positions together with the neutrino best-fit positions and both contours in Fig 16. We also report a coincidence between the ’prototype’ BFF AT2017bgt and the high-energy neutrino event IC-200410A. This event is, however, a poorly reconstructed cascade event, therefore we do not discuss it further here.
The neutrino associated with AT2021loi was detected days and days after the first and second optical peaks, respectively. Notably, all three nuclear transients (including AT2019aalc) associated with neutrinos exhibited a similar pattern, with neutrino detection occurring days after the optical peaks. This might be explained with delayed outflow activity (van Velzen et al. 2024). In the case of AT2021loi, the inverted radio spectrum (Makrygianni et al. 2023) is compatible with a newly ejected outflow component, similarly to the case of AT2019aalc. Interactions between the out-flowing material and clouds in the surrounding of the SMBH (Wu et al. 2022) or between ultra-high energy protons (accelerated in a jet or outflow) and infrared photons from the dust echo (Winter & Lunardini 2023) were proposed to explain the observed neutrino emission of the nuclear transients. These findings indicate that outflows play an important role in the neutrino production. A case study for AT2021loi would help us to further constrain the neutrino production models of nuclear transient sources.
If the connection between the high-ionization coronal line detection and the unusually large dust echos of nuclear transients is real, this can be used to reveal more sources with strong dust echo and therefore more candidate neutrino sources. Notably, both AT2019aalc and AT2021loi are characterized by significant dust echos (van Velzen et al. 2024; Hinkle 2024: see also Subsect. 4.2) and unusually strong high-ionization coronal lines (Makrygianni et al. 2023: see also Subsect. 5.1).
6 Conclusions
We conducted a multi-wavelength campaign during the re-brightening of the nuclear transient AT2019aalc associated with a high-energy neutrino event. We found that the second optical flare is more luminous and decays slower than the first one. Both flares evolve rapidly and with constant color. The UV emission peaks at extremely high luminosities (at erg/s). The re-brightening is accompanied by an IR flare, similarly to the first flare when it was explained with dust echo emission. An X-ray flare peaked before the optical/UV emission, preceding two subsequent flares that were nearly an order of magnitude brighter and predominantly blackbody in nature. The X-ray flares are very soft which explains the unusually strong He ii lines. In the optical spectra, persistent Bowen fluorescence lines were detected over a year after the second flare. With VLBI we clearly detect a non-thermal radio source which favors an AGN-like origin. The radio flux monitoring revealed a long-term flare and an inverted spectrum above GHz, which is likely explained by a newly ejected, self synchrotron-absorbed outflow component. Altogether, our results suggest a Bowen Fluorescence Flare classification of AT2019aalc. We also classify AT2019aalc as an Extreme Coronal Line Emitter.
The two distinct optical flares might be attributed to the partial disruption of a single star within an AGN environment. The stellar debris-disk interaction in such a system could result in complex light curve variations e.g., the plateau observed in optical/UV (Chan et al. 2020). A TDE-related scenario provides an explanation not only for the suddenly enhanced activity of the nucleus but also for the high metallicity of the disk gas which was notably suggested by TDE-AGN simulations (Ryu et al. 2024). Future observations will help us to better understand the nature of AT2019aalc: a repeating TDE would lead to a roughly periodic flaring, while some other AGN activity would lead to erratic future flares.
Motivated by the high-energy neutrino associated with AT2019aalc, we cross-matched the ICECAT-1 with the positions of the known 4 BFFs. We found that AT2021loi is coincident with the high-energy neutrino event IC230511 confirming the flaring AGN-neutrino connection studied by van Velzen et al. (2024). Apart from the known multi-wavelength properties of BFFs, AT2021loi shares characteristics with AT2019aalc such as late-time IR flare (due to dust echo), high-ionization coronal lines in its spectra (and also -band excess), radio detection (and inverted radio spectrum) and re-brightening episode in optical. In both cases the AGN are classified as broad-line AGN in-prior the transients occurred.
Acknowledgements.
We thank Benny Trakhtenbrot for useful discussions. PMV, AF, BA, ST, AK, DJB, EM-B acknowledge the support from the DFG via the Collaborative Research Center SFB1491 Cosmic Interacting Matters - From Source to Signal. EH acknowledges support by NASA under award number 80GSFC21M0002.The European VLBI Network is a joint facility of independent European, African, Asian, and North American radio astronomy institutes. Scientific results from data presented in this publication are derived from the following EVN project code: EV027. e-MERLIN is a National Facility operated by the University of Manchester at Jodrell Bank Observatory on behalf of STFC. We acknowledge the use of public data from the Swift data archive. This research has made use of the XRT Data Analysis Software (XRTDAS) developed under the responsibility of the ASI Science Data Center (ASDC), Italy. The Australia Telescope Compact Array is part of the Australia Telescope National Facility (https://ror.org/05qajvd42) which is funded by the Australian Government for operation as a National Facility managed by CSIRO. We acknowledge the Gomeroi people as the Traditional Owners of the Observatory site. Funding for the Sloan Digital Sky Survey V has been provided by the Alfred P. Sloan Foundation, the Heising-Simons Foundation, the National Science Foundation, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. SDSS telescopes are located at Apache Point Observatory, funded by the Astrophysical Research Consortium and operated by New Mexico State University, and at Las Campanas Observatory, operated by the Carnegie Institution for Science. The SDSS web site is www.sdss.org. SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration, including Caltech, The Carnegie Institution for Science, Chilean National Time Allocation Committee (CNTAC) ratified researchers, The Flatiron Institute, the Gotham Participation Group, Harvard University, Heidelberg University, The Johns Hopkins University, L’Ecole polytechnique fédérale de Lausanne (EPFL), Leibniz-Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Extraterrestrische Physik (MPE), Nanjing University, National Astronomical Observatories of China (NAOC), New Mexico State University, The Ohio State University, Pennsylvania State University, Smithsonian Astrophysical Observatory, Space Telescope Science Institute (STScI), the Stellar Astrophysics Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Illinois at Urbana-Champaign, University of Toronto, University of Utah, University of Virginia, Yale University, and Yunnan University. This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, and NEOWISE, which is a project of the Jet Propulsion Laboratory/California Institute of Technology. WISE and NEOWISE are funded by the National Aeronautics and Space Administration. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This work has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The Asteroid Terrestrial-impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile. This work is based on data from eROSITA, the soft X-ray instrument aboard SRG, a joint Russian-German science mission supported by the Russian Space Agency (Roskosmos), in the interests of the Russian Academy of Sciences represented by its Space Research Institute (IKI), and the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The SRG spacecraft was built by Lavochkin Association (NPOL) and its subcontractors, and is operated by NPOL with support from the Max Planck Institute for Extraterrestrial Physics (MPE). The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg & ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. These results made use of the Lowell Discovery Telescope (LDT) at Lowell Observatory. Lowell is a private, non-profit institution dedicated to astrophysical research and public appreciation of astronomy and operates the LDT in partnership with Boston University, the University of Maryland, the University of Toledo, Northern Arizona University and Yale University. The upgrade of the DeVeny optical spectrograph has been funded by a generous grant from John and Ginger Giovale and by a grant from the Mt. Cuba Astronomical Foundation. Some of the data presented herein were obtained at Keck Observatory, which is a private 501(c)3 non-profit organization operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the Native Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain. This publication makes use of data products from the Two Micron All Sky Survey, which is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center/California Institute of Technology, funded by the National Aeronautics and Space Administration and the National Science Foundation. Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-2034437 and a collaboration including Caltech, IPAC, the Weizmann Institute for Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, and IN2P3, France. Operations are conducted by COO, IPAC, and UW.
Appendix A Spetroscopic details
Date | Telescope | Instrument | Exp. time | Slit width | Seeing FWHM |
(s) | (arcsec) | (arcsec) | |||
Post-flare-1 | |||||
2021-07-06 | Keck-I | LRIS | 300 | 1.0 | 1.1 |
Over-flare-2 | |||||
2023-06-15 | Keck-I | LRIS | 300 | 1.0 | 1.4 |
2023-06-22 | LDT | DeVeny | 800 | 1.5 | 1.2 |
2023-07-11 | Palomar Hale m | DBSP | 180 | 1.5 | 1.5 |
2023-08-17 | Palomar Hale m | DBSP | 900 | 1.5 | 1.1 |
2024-05-21 | Palomar Hale m | DBSP | 1200 | 1.5 | 1.3 |
Date | Instrument | Resolving power | |||
(approx.) | (days) | ||||
2008-04-05 | SDSS | 1500-2500 | -11 years | 2a,b | |
2021-06-07 | LRIS | 1000 | -737 | 0.21 0.01 | 4a,b,c,d |
2023-06-15 | LRIS | 1000 | -28 | 0.64 0.02 | 4a,b,c,d |
2023-06-22 | Deveny | 750 | -21 | 0.50 0.01 | 4a,b,c,d |
2023-07-11 | DBSP | 1000 | -2 | 0.83 0.04 | 4a,b,c,d |
2023-08-17 | DBSP | 1000 | +35 | 0.25 0.01 | 5a,b,c,d,e |
2024-05-21 | DBSP | 1000 | +315 | 0.30 0.02 | 5a,b,c,d,e |